OpenStudy (blondisaurus):

If c(x)= 4x^3-5x^2+2 and d(x)= 3x^2+ 6x-10, find each value Equation: 6c(4a)+2d(3a-5)

OpenStudy (phi):

6 c(4a) is short hand for replace x with (4a) in the expression 4x^3-5x^2+2 (everywhere you see x, put in (4a) instead) then multiply every term by 6 that gives you 6 c(4a) use the same idea for 2 d(3a-5) where d(x) is 3x^2+ 6x-10 Can you do that?

OpenStudy (blondisaurus):

I did it the way you said but didnt come up to the answer in the book

OpenStudy (phi):

not too surprising, it looks messy. what did you get for just c(4a) ?

OpenStudy (blondisaurus):

96a^3-120a^2+12

OpenStudy (phi):

that is not c(4a)

OpenStudy (phi):

replace x with (4a) in the expression 4x^3-5x^2+2

OpenStudy (blondisaurus):

I did and i got 16a^3-20a^2+2. Then i multiplied everything by 6 like you said :o

OpenStudy (phi):

ok, first when you replace x with 4a in 4* x^3 you write it like this: 4* (4a)^3 remember that x^3 is short-hand for x*x*x (x times itself 3 times) so the answer with 4a instead of x is 4* (4a)*(4a)*(4a) you do not get 16a^3 can you retry?

OpenStudy (blondisaurus):

256^3-80a^2+2?

OpenStudy (phi):

you would reorder 4* (4a)*(4a)*(4a) = 4*4*a*4*a*4*a as 4*4*4*4 * a*a*a or using short-hand (exponents) 4^4 * a^3 or 256 a^3

OpenStudy (phi):

256^3-80a^2+2? yes, assuming you left out an a in there. now multiply by 6 6(256a^3 -80a^2 +2)

OpenStudy (phi):

I get 1536 a^3 -480 a^2 +12 now the ugly one d(3a-5) where 3x^2+ 6x-10 if we just replace x with (3a-5) we get 3*(3a-5)^2 +6(3a-5) -10 can you simplify this?

OpenStudy (blondisaurus):

27a^2+18a-115?

OpenStudy (phi):

27 a^2 -72a +35

OpenStudy (phi):

what did you get for (3a-5)^2 ? (3a-5)(3a-5)= 9a^2 -15a -15a +25 9a^2 -30a+25 we want 3*(3a-5)^2 so mult by 3: 27a^2 -90a +75

OpenStudy (phi):

can you finish by adding in +6(3a-5) -10

OpenStudy (blondisaurus):